# Price Experimentation and Interference

### v2

@@ 8 Conclusion and Future Work
In this paper we studied biases that arise due to interference in platform experiments, with a particular emphasis on profit or revenue metrics. Notably, we find that in certain settings, these biases can be so severe that they could introduce a “change of sign” in estimation of profit or revenue.
In future work, we aim to extend the analysis to other settings of interest, including two-sided market settings where the variable of interest is the percentage fee charged by the platform (as a fraction of gross revenue). Further, our findings in Section 7 motivate further empirical study of the prevalence of changes of sign in practice. It is natural to consider more robust empirical calibration strategies, as well as practical methodological approaches to mitigating this change of sign. These remain important directions for future work.

### v3

@@ 9 Conclusion and Future Work
In this paper we studied biases that arise due to interference in platform experiments, with a particular emphasis on profit or revenue metrics. Notably, we find that in certain settings, these biases can be so severe that they could introduce a “change of sign” in estimation of profit or revenue. We then show a novel method for debiasing canonical A/B experiments, by splitting half of subjects into treatment and half into control.
In future work we aim to extend our debiasing result to allow for correcting canonical experiments beyond the equal split between control and treatment. We also aim to run empirical studies to get a better sense of the prevalence of changes of sign in practice.